OpenStudy (anonymous):

Solve the rational equation: 2-4/x^2=-2/x

OpenStudy (anonymous):

Try the method of cross-multiplication first.\[\huge \frac{a}{b}=\frac{c}{d}\ \Rightarrow \ ad=bc\]

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

is the equation \[\Large 2 - \frac{4}{x^2}=-\frac{2}{x}\] ?

OpenStudy (anonymous):

Yes =)

jimthompson5910 (jim_thompson5910):

alright great, the first thing we do is get rid of the fractions

jimthompson5910 (jim_thompson5910):

we do this by multiplying every term by the LCD x^2

jimthompson5910 (jim_thompson5910):

So multiply that first term 2 by x^2 to get 2x^2 Then multiply -4/x^2 by x^2 to get x^2*(-4/x^2) = -4 Finally, multiply -2/x by x^2 to get x^2(-2/x) = -2x^2/x = -2x So 2-4/x^2=-2/x turns into 2x^2 - 4 = -2x

jimthompson5910 (jim_thompson5910):

What's next from here?

OpenStudy (anonymous):

Then would I get: 4x-4=-2x

jimthompson5910 (jim_thompson5910):

No, 2x^2 is not the same as 4x

OpenStudy (anonymous):

Okay, can you explain how I could get x alone?

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

First get the expression into standard form 2x^2 - 4 = -2x 2x^2 - 4 + 2x = 0 2x^2 + 2x - 4 = 0 Now use the quadratic formula to solve for x. x = (-b+-sqrt(b^2-4ac))/(2a) x = (-(2)+-sqrt((2)^2-4(2)(-4)))/(2(2)) x = (-2+-sqrt(4-(-32)))/(4) x = (-2+-sqrt(36))/4 x = (-2+sqrt(36))/4 or x = (-2-sqrt(36))/4 x = (-2+6)/4 or x = (-2-6)/4 x = 4/4 or x = -8/4 x = 1 or x = -2 So the solutions are x = 1 or x = -2

OpenStudy (anonymous):

Thank you!! =)

jimthompson5910 (jim_thompson5910):

yw